# Density of states and non-smooth Lyapunov exponent in the localized phase

### v1

@@ Abstract
Localization of wave functions in the disordered models can be characterized by the Lyapunov exponent, which is zero in the extended phase and nonzero in the localized phase. Previous studies have shown that this exponent is a smooth function of eigenenergy in the same phase, thus its non-smoothness can serve as strong evidence to determine the phase transition from the extended phase to the localized phase. However, logically, there is no fundamental reason that prohibits this Lyapunov exponent from being non-smooth in the localized phase. In this work, we show that if the localization centers are inhomogeneous in the whole chain and if the system possesses (at least) two different localization modes, the Lyapunov exponent can become non-smooth in the localized phase at the boundaries between the different localization modes. We demonstrate these results using several slowly varying models and show that the singularities of density of states are essential to these non-smoothness, according to the Thouless formula. These results can be generalized to higher-dimensional models, suggesting the possible delicate structures in the localized phase, which can revise our understanding of localization hence greatly advance our comprehension of Anderson localization.
@@
Anderson localization is one of the most fundamental concepts in condensed matter physics and quantum simulation [1, 2, 3, 4, 5, 6, 7, 8, 9]. Recently, significant attention has been focused on this transition in various disordered and quasiperiodic systems, which can be realized in experiments using cold atoms [10, 11, 12, 13, 14, 15, 16], optical systems [17, 18, 19, 20, 21], and superconducting qubits [22, 23, 24]. Furthermore, their Lyapunov exponent $\gamma(E)$ and the associated mobility edges [25, 26, 27, 28], which characterize the Anderson phase transition, can be calculated using transfer matrix method [29, 30, 31], dual transformation method [32, 33, 34, 35, 36] and Avila’s global theory [37, 38, 39]. In these studies, the Lyapunov exponent is presumed to be a smooth function of energy $E$ within the same phase, and the non-smooth transition in $\gamma(E)$ from zero (for extended phase) to non-zero (for localized phase) has been regarded as smoking-gun evidence of phase transition [40, 41, 42]. However, with this wisdom, a fundamental question remains unexplored: can the Lyapunov exponent exhibit non-smooth behavior within a single localized phase? At first glance, such behavior appears to contradict the standard picture, where non-smoothness of $\gamma(E)$ signals a boundary between distinct phases [38, 43]. Yet, there is no fundamental principle to forbid the Lyapunov exponent from being non-smooth within the localized phase, with $\gamma(E)>0$. This suggests the possibility of much more subtle structures inside the localized phase than previously appreciated, which will greatly enhance our comprehension of Anderson localization in disordered models.
@@
In this manuscript, we demonstrate that (I) the Lyapunov exponent $\gamma(E)$ can exhibit non-smooth behavior within the localized phase, without indicating a transition to an extended phase, and (II) this occurs when the system possesses multiple localization modes. Specifically, we study a class of slowly varying models which supports the localization around either potential nodes or extrema, naturally giving rise to two localization modes. At the boundaries separating these modes, we observe non-smooth behavior in $\gamma(E)$. A more detailed consideration of the analysis based on the Thouless formula [44] shows that the divergences in DOS play a crucial role in the non-smooth transition between localized states. Our results suggest that the localized phase may possess much more delicate structures that warrant more careful consideration in future investigation, which could significantly advance our understanding of localization phenomena.
@@
Non-smooth Lyapunov exponent in the localized phase.—We begin by analyzing the localization properties in disordered systems. In these models, the on-site potentials can be regarded as homogeneous in the sense of ergodicity, thus the localization centers of the wave functions will be uniformly distributed in the whole chain. Consequently, these models exhibit only a single localization mode, implying that the Lyapunov exponent $\gamma(E)$ should be a smooth function of $E$. Here, $E$ is not necessary to be the eigenenergy of the model. This idea further suggests that for $\gamma(E)$ to be a non-smooth function of $E$ in the localized phase, the system must possess at least two distinct localization modes, such as different localization centers. We implement this concept through the following slowly varying potential model [45, 46, 47, 48, 49, 50, 51]
@@
where $c_{i}^{\dagger}$ ($c_{i}$) denotes the creation (annihilate) operator at site $i$, $t$ represents the nearest-neighbor hopping term, and $L$ is the chain length. We first adopt the potential as $V_{i}=V\cos(\pi\alpha i^{\nu})$ with $10<\nu<$ 1 [45, 46, 47, 48, 49, 50, 51], which supports two distinct l shown in Fig. 1(a), that is, the wave functions are likely to be localized either at potential nodes or potential extrema when $V$ is sufficiently large. In this way, the localization centers (mean position of wave function) exhibit inhomogeneity.
@@
The form of $V_{i}$ allows for the constant approximations in the large-$i$ limit, facilitating analytical treatment of localization properties. In the bare hopping regime ($V=0$), the kinetic energy is bounded by $|K|\leq 2t$ (with $t>0$). Consequently, when the wave functions are localized at the potential extrema, their energy should be either $K-V$ or $K+V$, defining the boundary curves ±$\pm 2t\pm V$, which intersect at $V=2t$. In contrast, when states are fully localized at the potential nodes, their energy is expected to be independent of the potential depth, leading to $E\approx 0$. It will depends on $V$ when the wave functions are spatially localized near the nodes. This argument yields the phase diagram shown in Fig. 1(b), which has been rigorously established in Ref. [48]. The key insight is that the presence of two distinct localization modes allows the localization length and the associated Lyapunov exponent to exhibit non-smooth variations as functions of energy $E$ in the localized phase. More specifically, if we denote the Lyapunov exponents in these two cases as $\gamma_{\text{nl}}(E)$ and $\gamma_{\text{el}}(E)$, the non-smooth behavior occurs at $\gamma_{\text{nl}}(E)=\gamma_{\text{el}}(E)$, as illustrated in Fig. 1(c). It should be emphasized that these states belong to the same phase in the sense that their measured properties, such as conductance will be unchanged in the thermodynamic limit. From a much broader perspective, this phenomenon highlights that even within a single phase characterized by specific order parameters, certain physical properties may still exhibit some non-smooth behaviors.
@@
$\gamma(E)$, DOS, and Thouless formula.—We next perform an analytical investigation of the underlying physics and subsequently validate our results through numerical simulations. The Lyapunov exponent can be obtained from the Thouless formula [44, 52, 53, 54]
[…]
where the DOS is defined as $D(E)=\sum_{n=1}^{L}\delta(E-E_{n})$. This relation, termed as Hilbert transformation, implies that singularities in $D(E)$ can yield non-smooth behavior in $\gamma(E)$. This behavior can be categorized into two distinct cases, one with $\gamma(E)$ changes from zero to a finite value corresponding to Anderson transition [46], and the other with $\gamma(E)$ varies between two finite values representing transition between different localization modes. For the Aubry-André-Harper (AAH) model [55, 56, 57, 58], $\gamma(E)$ remains zero for all energies $E$ when $V<2t$, while for $|V|>2t$, it is given by $\gamma=\text{ln}(|V/2t|)$, indicating a phase transition at $V=\pm 2t$, as determined by self-duality [55]. Consequently, although $\gamma(E)$ remains continuous as a function of $E$, its first derivative $\partial\gamma$ exhibits discontinuities at the phase boundaries, signaling a phase transition. This constitutes a major result of this work, that is, the Lyapunov exponent can be a continuous yet non-smooth function of $E$ in the localized phase , which always indicates a transition between different kinds of states.
@@
For the slowly varying potential $V_{i}=V\cos(\pi\alpha i^{\nu})$, the numerical results for $\gamma(E)$ and its first derivative $\partial\gamma$ at $V=t$ and $V=3t$ are presented in Fig. 2. When $V=t$, the DOS diverges at $E=\pm t$ [see Fig. 2(a)], with the Lyapunov exponent remaining zero for $|E|<t$, characterizing an extended phase [see Fig. 2(b)]. For $|E|>t$, the Lyapunov exponent approximately follows $\gamma\propto|E|-t$ around the critical points $E_{\rm c}=\pm t$. These results indicate that the phase transition, characterized by non-smooth behavior of $\gamma$, is driven by the divergence of DOS. When $V=3t$, all eigenstates are localized with $\gamma>0$ [see the black solid line in Fig. 2(d)], and the DOS still diverges at $E=\pm t$ [see the black solid line in Fig. 2(b)]. Meanwhile, we still observe non-smoothness of $\partial\gamma$ in this case. All of those evidence indicate the existence of two different localized states.
@@
To gain deeper insight into the nature of different localization modes, we calculate the localization center as $\bar{i}=\sum_{i}i|\psi_{i}|^{2}$, where $\psi$ denotes a normalized eigenstate. For node localized states, we expect $V_{\bar{i}}\sim 0$, whereas for extreme localized states, we expect $V_{\bar{i}}\sim\pm V$. Focusing on the regime $V>2t$, we set $V=3t$ and plot the potential at the localization center as a function of $E$ in Fig. 3(a). In this case, the boundaries between these localized states occur at $E=\pm t$, which agree well with our expectations. For extreme localized states, we find $V_{\bar{i}}\rightarrow\pm V$, while for node localized states, we observe $|V_{\bar{i}}|<V$. Three typical wave functions corresponding to these localization modes are presented in Fig. 3(b). These results reveal that the boundaries between two localization modes give rise to the non-smoothness of $\gamma(E)$, providing a physical interpretation of this phenomenon from the perspective of wave function.
@@
The above results may be found in much more general slowly varying potentials. To this end, we have also verified that when the on-site potential is given by [33, 59, 60]
@@
with $|b|<1$ and $10<\nu<$ 1, the same non-smoothness of $\gamma(E)$ can be obtained. Using the same approach, we find that the phase boundaries can be described by $E=\pm 2t+V/(1-b)$ and $E=\pm 2t-V/(1+b)$. In this way, all states are localized when $V>V_{c}=2t(1-b^{2})$, and the boundaries between node and extreme localized states occur at $E=-2t+V/(1-b)$ and $E=2t-V/(1+b)$ when $V>V_{c}$. The singularities in DOS lead to the non-smooth behavior of the Lyapunov exponents in the localized phase. In Ref. [61], it has been found that this model support critical phase when $|b|>1$ and $\nu=1$, and the condition with $\nu<1$ will be considered elsewhere.
Figure 4: (a) Fractal dimension $D_{2}$ versus eigenenergy $E$ and hopping strength $\lambda$ for the quasiperiodic slowly varying hopping model described by Eq. (5), with $b=0.3$. (b) Three representative eigenstates as marked in Fig. 4(a). (c)-(d) DOS, and (e)-(f) Lyapunov exponent $\gamma$ and its derivative $\partial\gamma$ versus $E$ for different $\lambda$. The dashed lines in (c) and (e) represent the boundaries $E=7t/13$ and $E=19t/13$, whereas in (d) and (f), the boundaries are located at $E=-21t/13$ and $E=47t/13$.
Slowly varying hopping model.— Furthermore, we consider the model with slowly varying hopping, described by the Hamiltonian [49]
@@
where $t$ is a constant, and $\lambda$ characterizes the amplitude of the slowly varying hopping modulation. When $|\lambda|<(1+b)t$, all sites are connected, allowing the existence of extend states. However, when $|\lambda|>(1+b)t$, there exist infinitely many indices $i$ satisfying $|g_{\nu}(i)|<\varepsilon$, where $\varepsilon$ is an infinitesimally small positive constant. Such hopping terms divide the entire system into weakly coupled blocks, leading to spatially localized states. This system also supports two different localization modes: node localization at $g_{\nu}(i)\sim 0$ and extreme localization at $g_{\nu}(i)\sim t+\lambda/(1-b)$. Following the similar approach [48, 49, 51], we derive the boundaries for different localization modes at
@@
The phase diagram obtained from numerical simulations is presented in Fig. 4(a). In Fig. 4(b), we demonstrate that these states are indeed spatially localized at distinct localization centers. Meanwhile, we calculate the DOS for $V=t$ and $V=3t$, as shown in Figs. 4(c) and 4(d). From the DOS, we can determine the Lyapunov exponent $\gamma(E)$ and its first derivative $\partial\gamma$, which are presented in Figs. 4(e) and 4(f). For $V=t$, we find that $\gamma(E)=0$ when $137t/13<E<19t/$ 13, indicating a localized-delocalized transition driven by the divergence of DOS. When $V=3t$, $\gamma(E)$ remains positive, indicating the localization of all states. We finds discontinuous DOS and non-smooth Lyapunov exponents at $E=-21t/13$ and $E=47t/13$.
@@
Absence of node localized states.—Building on the above examples and insights, we arrive at the central conclusion of this study. In following, we provide a counterexample to further substantiate this conclusion. Specifically, we consider a model characterized by the following slowly varying square-well potential [46, 48]
@@
where $\text{sign}(x)$ denotes the sign function. Since this potential lacks nodes, it does not support node localization, yielding a single localization mode. To verify this, we plot the fractal dimension $D_{2}$ versus energy $E$ and potential strength $V$ in Fig. 5(a), with phase boundaries determined by $E=\pm(2t\pm V)$. We confirm that in the overlapped regime for these four curves, the wave functions are extended, whereas in the un-overlapped regime, wave functions are localized at the potential extrema as illustrated in Figs. 5(c) and 5(d). Additionally, we calculate the DOS and the corresponding Lyapunov exponent, demonstrating that both $\gamma(E)$ and its first-order derivative are smooth functions of $E$, as presented in Fig. 5(b). Since this model only supports a single localization mode, the non-smooth behavior in $\gamma(E)$ is absent, leading to a smooth Lyapunov exponent similar to those observed in disordered and quasiperiodic potential models [29].

### v2

# Divergent density of states and non-analytic Lyapunov exponent in one-dimensional slowly varying systems

@@ Abstract
Localization of wave functions in disordered systems can be characterized by the Lyapunov exponent, which is zero in the extended phase and nonzero in the localized phase. Previous studies have shown that this exponent is an analytic function of eigenenergy in a given phase, thus its non-analytic behavior has been commonly used to determine the boundaries between the extended and localized phases. In this work, we show that if the localization centers are inhomogeneous across the whole chain and the system possesses (at least) two different localization modes, the Lyapunov exponent can become non-analytic in the localized phase at the boundaries between the different localization modes. We establish this central result by using several one-dimensional slowly varying models, and reveal that the non-analytic feature in the Lyapunov exponent is inherently tied to the singularities in the density of states through the Thouless formula. The possible existence of delicate structures in the localized phase effectively broadens our understanding of Anderson localization.
I Introduction
Anderson localization is one of the most fundamental concepts in condensed matter physics, which describes the suppression of wave transport caused by disorder-induced interference [1, 2, 3, 4, 5, 6, 7, 8, 9]. Over the past decades, rapid experimental progress has enabled the controlled realization of disordered and quasiperiodic systems, allowing the localization–delocalization transition to be observed. These transitions have been experimentally observed in a broad range of experimental platforms, including cold atoms [10, 11, 12, 13, 14], optical systems [15, 16, 17], and superconducting qubits [18, 19]. These experimental realizations have greatly enriched our understanding of how disorder shapes localization phenomena.
@@ I Introduction
A central theoretical tool for describing Anderson localization is the Lyapunov exponent $\gamma(E)$, which characterizes the exponential decay of wave functions and thereby distinguishes extended states ($\gamma=0$) from localized ones ($\gamma>0$) [20, 21]. The energy dependence of $\gamma(E)$ encodes key information about the system’s localization properties, serving as a powerful indicator of mobility edges (MEs) and phase boundaries. Several complementary methods have been developed to obtain $\gamma(E)$ and the associated MEs, such as the transfer-matrix approach [22, 23, 24], dual transformation method [25, 26, 27, 28, 29] and Avila’s global theory [30, 31, 32]. In these studies, the Lyapunov exponent is assumed to be an analytic function of energy $E$ within the same phase, and the non-analytic transition in $\gamma(E)$ from zero to non-zero has been regarded as smoking-gun evidence of phase transition [33, 34, 35, 36, 37]. However, with this wisdom, a fundamental question remains unexplored: can the Lyapunov exponent exhibit non-analytic behavior within a localized phase? At first glance, such behavior appears to contradict the standard picture, where non-analyticity of $\gamma(E)$ signals a boundary between distinct phases [31]. Yet, there is no fundamental principle to forbid the non-analytic Lyapunov exponent in the localized phase, with $\gamma(E)>0$. This suggests the possibility of much more subtle structures inside the localized phase than previously appreciated, which will greatly enhance our comprehension of Anderson localization in disordered models.
@@
In this manuscript, we explore this intriguing question by analyzing a class of one-dimensional slowly varying models that host multiple localization modes. We demonstrate that (I) the Lyapunov exponent $\gamma(E)$ can exhibit non-analytic behavior within the localized phase, without indicating a transition to an extended phase, and (II) this occurs when the system possesses multiple localization modes. We first show that a slowly varying potential which supports the localization around either potential nodes or extrema naturally gives rise to two localization modes. We then prove the emergence of non-analytic behavior in $\gamma(E)$ at the boundaries separating these localization modes. A detailed analysis based on the Thouless formula [38] shows that the divergences in the density of states (DOS) play a crucial role in the non-analytic transition between localized states. Our results suggest that the localized phase may possess much more delicate structures than previously recognized, warranting more careful consideration in future studies and offering significant insight into our understanding of localization phenomena.
The rest of this manuscript is organized as follows. Section II takes slowly varying potentials as a representative example to illustrate that the non-analytic Lyapunov exponent arises from the coexistence of two kinds of localized modes in a localized phase. Section III extends this analysis to more general slowly varying models and presents additional supporting evidence. Finally, Sec. IV summarizes and discusses our results.
@@
II Slowly varying potential model
II.1 Dual localization modes
We begin by analyzing the localization properties of conventional disordered systems, which is helpful for understanding the non-analytic behavior of the Lyapunov exponent. In these models, the on-site potentials can be regarded as homogeneous in the sense of ergodicity, thus the localization centers of the wave functions will be uniformly distributed in the whole chain. Consequently, these models exhibit only a single localization mode, implying that the Lyapunov exponent $\gamma(E)$ should be an analytic function of $E$ [31]. This idea further suggests that for $\gamma(E)$ to be a non-analytic function of $E$ in the localized phase, the system must possess at least two distinct localization modes, such as different localization centers. We implement this concept through the following slowly varying potential model [39, 40, 41, 42, 43, 44, 45]
@@ II.1 Dual localization modes
where $c_{i}^{\dagger}$ ($c_{i}$) denotes the creation (annihilation) operator at site $i$, $t$ represents the nearest-neighbor hopping term, and $L$ is the chain length. If the potential is taken as $V_{i}=V\cos(\pi\alpha i^{\nu})$ with $0<\nu<1$ [39, 40, 41, 42, 43, 44, 45], there exist two distinct localization modes as shown in Fig. 1(a), that is, the wave functions are likely to be localized either at potential nodes or potential extrema when $V$ is sufficiently large. In this way, the localization centers (mean position of wave function) exhibit spatial inhomogeneity, marking the emergence of dual localization centers in the system. The inhomogeneity of the localization centers is the reason for the non-analytic behavior of Lyapunov exponent.
@@
The form of $V_{i}$ allows for the constant approximations in the large-$i$ limit, facilitating analytic treatment of localization properties. In the bare hopping regime ($V=0$), the kinetic energy is bounded by $|K|\leq 2t$ (with $t>0$). Consequently, when the wave functions are localized at the potential extrema, their energy should be either KVK-V or $K+V$, defining the boundary curves $\pm 2t\pm V$, which intersect at $V=2t$. In contrast, when states are fully localized at the potential nodes, their energy is expected to be independent of the potential depth, leading to $E\approx 0$. It will depend on $V$ when the wave functions are spatially localized near the nodes. This argument yields the phase diagram shown in Fig. 1(b). While the boundaries between the extended and localized states seen in Fig. 1(b) have been rigorously established in Ref. [42], the key insight here is that the presence of two distinct localization modes allows the localization length and the associated Lyapunov exponent to exhibit non-analytic variations as functions of energy $E$ in the localized phase. More specifically, if we denote the Lyapunov exponents in the two localization modes as $\gamma_{\text{nl}}(E)$ and $\gamma_{\text{el}}(E)$, the non-analytic behavior is bound to occur at $\gamma_{\text{nl}}(E)=\gamma_{\text{el}}(E)$, as illustrated in Fig. 1(c). It should be emphasized that these states belong to the same phase in the sense that their measured properties, such as conductance will be unchanged in the thermodynamic limit. From a broader perspective, this phenomenon highlights that even within a single phase characterized by specific order parameters, certain physical properties may still exhibit some non-analytic behaviors.
II.2 Lyapunov exponent, DOS, and Thouless formula
According to the Thouless formula, the Lyapunov exponent can be obtained by [38, 46]
$\gamma(E)=\int dE^{\prime}D(E^{\prime})\ln|E-E^{\prime}|,$ (2)
where the DOS is defined as
$D(E)=\sum_{n=1}^{L}\delta(E-E_{n}).$ (3)
This relation, termed as Hilbert transformation, implies that singularities in $D(E)$ can yield non-analytic behavior in $\gamma(E)$. This behavior can be categorized into two distinct cases, one with $\gamma(E)$ changing from zero to a finite value corresponding to Anderson transition [40], and the other with $\gamma(E)$ varying between two finite values representing transition between different localization modes. For the Aubry-André-Harper (AAH) model [47, 48, 49, 50], $\gamma(E)$ remains zero for all energies $E$ when $V<2t$, while for $|V|>2t$, it is given by $\gamma=\text{ln}(|V/2t|)$, indicating a phase transition at $V=\pm 2t$, as determined by self-duality [47]. Consequently, although $\gamma(E)$ remains continuous as a function of $E$, its first derivative $\partial\gamma$ exhibits discontinuities at the phase boundaries, signaling a phase transition. This constitutes a major result of this work, that is, the Lyapunov exponent can be a continuous yet non-analytic function of $E$ in the localized phase, which always indicates a transition between different kinds of states.
II.3 Methods for characterizing localization properties
Here, we summarize the numerical methods used to characterize the localization properties of the system, including the fractal dimension and the Lyapunov exponent.
II.3.1 Fractal dimension $D_{2}$
In the theory of Anderson localization, the inverse participation ratio (IPR) is a standard measure to quantify the localization properties of eigenstates, defined as
${\rm IPR}^{n}=\sum_{i=1}^{N}|\psi_{i}^{n}|^{4},$ (4)
Here, $\psi_{i}^{n}$ is the normalized wavefunction at site $i$, with $n$ being the band index. In the limiting cases, ${\rm IPR}^{n}=1$ for a state localized on a single site ($L^{0}$ scaling), while ${\rm IPR}^{n}\propto L^{-1}$ for a uniformly extended state. Based on this, we can define the fractal dimension as
$D_{2}^{n}=-\lim_{L\rightarrow\infty}\frac{{\rm ln}{\rm IPR}^{n}}{{\rm ln}L},$ (5)
which falls between 0 and $1$. Here, $D_{2}$ tends to 1 for extended states and 0 for localized states, while $0<D_{2}<1$ indicates multifractal states.
II.3.2 Lyapunov exponent $\gamma$
The Lyapunov exponent of a tight-binding model involving only nearest-neighbor hopping terms can be calculated as
$\gamma(E)=\lim\limits_{L\rightarrow\infty}\frac{1}{2L}\ln(\mathbf{T}_{L}^{\dagger}\mathbf{T}_{L}),\hskip 14.22636pt\mathbf{T}_{L}=\prod\limits_{i=1}^{N}T_{i},$ (6)
where the transfer matrix $T_{i}$ is
$T_{i}=\begin{pmatrix}{V_{i}-E}&-1\\ 1&0\end{pmatrix}$. (7)
Here, $\mathbf{T}_{L}$ is the product of single-site transfer matrices $T_{i}$ up to length $L$. In practice, the QR decomposition is employed to improve the numerical stability and efficiency in computing $\mathbf{T}_{L}$.
@@ II.3.2 Lyapunov exponent $\gamma$
II.4 Numerical results
@@
For the slowly varying potential $V_{i}=V\cos(\pi\alpha i^{\nu})$, the numerical results for the DOS, $\gamma(E)$ and its first derivative $\partial\gamma$ at $V=t$ and $V=3t$ are presented in Fig. 2. When $V=t$, the DOS diverges at $E=\pm t$ [see Fig. 2(a)], with the Lyapunov exponent remaining zero for $|E|<t$, characterizing an extended phase [see Fig. 2(b)]. For $|E|>t$, the Lyapunov exponent approximately follows $\gamma\propto|E|-t$ around the critical points $E_{\rm c}=\pm t$. These results indicate that the phase transition, characterized by non-analytic behavior of $\gamma$, is driven by the divergence of DOS. When $V=3t$, all eigenstates are localized with $\gamma>0$ [see the black solid line in Fig. 2(d)], and the DOS still diverges at $E=\pm t$ [see the black solid line in Fig. 2(b)]. Meanwhile, we still observe non-analyticity of $\partial\gamma$ in this case. All of the evidence indicates the existence of two different localized states.
@@
To gain deeper insight into the nature of different localization modes, we calculate the localization center as $\bar{i}=\sum_{i}i|\psi_{i}|^{2}$, where $\psi$ denotes a normalized eigenstate. For node localized states, we expect $V_{\bar{i}}\sim 0$, whereas for extreme localized states, we expect $V_{\bar{i}}\sim\pm V$. Focusing on the regime $V>2t$, we set $V=3t$ and plot the potential at the localization center as a function of $E$ in Fig. 3(a). In this case, the boundaries between these localized states occur at $E=\pm t$, which agree well with our expectations. For extreme localized states, we find $V_{\bar{i}}\rightarrow\pm V$, while for node localized states, we observe $|V_{\bar{i}}|<V$. Three typical wave functions corresponding to these localization modes are presented in Fig. 3(b). These results reveal that the boundaries between two localization modes give rise to the non-analyticity of $\gamma(E)$, providing a physical interpretation of this phenomenon from the perspective of wave function.
Figure 4: (a) Fractal dimension $D_{2}$ versus eigenenergy $E$ and hopping strength $\lambda$ for the slowly varying model described by Eq. (8), with $b=0.3$. (b)-(c) DOS, and (d)-(e) Lyapunov exponent $\gamma$ and its derivative $\partial\gamma$ versus $E$ for different $V$. The dashed lines in (b) and (d) represent the boundaries E=4t/7E=-4t/7 and $E=16t/13$, whereas in (d) and (f), the boundaries are located at E=4t/13E=-4t/13 and $E=16t/7$.
@@
The above results may be found in much more general slowly varying potentials. To this end, we have also verified that when the on-site potential is given by [26, 51, 52, 53]
@@
with $|b|<1$ and $0<\nu<1$, the same non-analyticity of $\gamma(E)$ can be obtained. Using the same approach, we find that the phase boundaries can be described by $E=\pm 2t+V/(1-b)$ and $E=\pm 2t-V/(1+b)$. In this way, all states are localized when $V>V_{c}=2t(1-b^{2})$, and the boundaries between node and extreme localized states occur at E=2t+V/(1b)E=-2t+V/(1-b) and E=2tV/(1+b)E=2t-V/(1+b) when $V>V_{c}$. The singularities in DOS lead to the non-analytic behavior of the Lyapunov exponents in the localized phase. The corresponding numerical results is presented in Fig. 4, which further demonstrate that the emergence of non-analytic $\gamma(E)$ is a general feature of slowly varying potentials supporting multiple localization modes.
III Slowly varying hopping model
To further generalize our analysis, we consider a model with slowly varying hopping, described by the Hamiltonian [43]
@@ III Slowly varying hopping model
where $t$ is a constant, and $\lambda$ characterizes the amplitude of the slowly varying hopping modulation. When $|\lambda|<(1+b)t$, all sites are connected, allowing the existence of extended states. However, when $|\lambda|>(1+b)t$, there exist infinitely many indices $i$ satisfying $|g_{\nu}(i)|<\varepsilon$, where $\varepsilon$ is an infinitesimally small positive constant. As a result, the hopping amplitudes vanish locally, dividing the chain into weakly coupled blocks and giving rise to spatially localized states. Similar to the slowly varying potential model, this system supports two distinct localization modes: node localization at $g_{\nu}(i)\sim 0$ and extreme localization at $g_{\nu}(i)\sim t+\lambda/(1-b)$. Following the similar approach [42, 43, 45], we derive the boundaries separating these localization modes as
@@
Figure 5: (a) Fractal dimension $D_{2}$ versus eigenenergy $E$ and hopping strength $\lambda$ for the quasiperiodic slowly varying hopping model described by Eq. (10), with $b=0.3$. (b) Three representative eigenstates as marked in Fig. 5(a). (c)-(d) DOS, and (e)-(f) Lyapunov exponent $\gamma$ and its derivative $\partial\gamma$ versus $E$ for different $\lambda$. The dashed lines in (c) and (e) represent the boundaries $E=7t/13$ and $E=19t/13$, whereas in (d) and (f), the boundaries are located at E=21t/13E=-21t/13 and $E=47t/13$.
The phase diagram obtained from numerical simulations is presented in Fig. 5(a). In Fig. 5(b), we demonstrate that these states are indeed spatially localized at distinct localization centers. Meanwhile, we calculate the DOS for $V=t$ and $V=3t$, as shown in Figs. 5(c) and 5(d). From the DOS, we can determine the Lyapunov exponent $\gamma(E)$ and its first derivative $\partial\gamma$, which are presented in Figs. 5(e) and 5(f). For $V=t$, we find that $\gamma(E)=0$ when $7t/13<E<19t/13$, indicating a localized-delocalized transition driven by the divergence of DOS. When $V=3t$, $\gamma(E)$ remains positive, indicating the localization of all states. We find discontinuous DOS and non-analytic Lyapunov exponents at E=21t/13E=-21t/13 and $E=47t/13$.
@@
Building on the above examples and insights, we arrive at the central conclusion of this study. In the following, we provide a counterexample to further substantiate this conclusion. Specifically, we consider a model characterized by the following slowly varying square-well potential [40, 42]
@@
where $\text{sign}(x)$ denotes the sign function. Since this potential lacks nodes, it does not support node localization, yielding a single localization mode. To verify this, we plot the fractal dimension $D_{2}$ versus energy $E$ and potential strength $V$ in Fig. 6(a), with phase boundaries determined by $E=\pm(2t\pm V)$. We confirm that in the overlapped regime for these four curves, the wave functions are extended, whereas in the non-overlapped regime, wave functions are localized at the potential extrema as illustrated in Figs. 6(c) and 6(d). Additionally, we calculate the DOS and the corresponding Lyapunov exponent, demonstrating that both $\gamma(E)$ and its first-order derivative are analytic functions of $E$, as presented in Fig. 6(b). Since this model only supports a single localization mode, the non-analytic behavior in $\gamma(E)$ is absent, leading to an analytic Lyapunov exponent similar to those observed in disordered and quasiperiodic potential models [22].
IV Conclusions